OpenStudy (katielong):

express the following in terms of 3 parital fractions: (2-5x^2) / (2-x)(1+x^2)

OpenStudy (katielong):

otherwise written as \[\frac{ 2-5x^2 }{ (2-x)(1+x)^2 }\]

OpenStudy (katielong):

so far, i have for the numerator: 2-5x^2 = A(1+x)(1+x)^2 + B(2-x)(1+x)^2 + C(2-x)(1+x)

OpenStudy (anonymous):

Your placement of the exponent is confusing. If \(1+x\) is being squared, you have \[\frac{2-5x^2}{(2-x)(1+x)^2}=\frac{A}{2-x}+\frac{B}{1+x}+\frac{C}{(1+x)^2}\] If you meant to have \(1+x^2\), then you have \[\frac{2-5x^2}{(2-x)(1+x^2)}=\frac{A}{2-x}+\frac{Bx+C}{1+x^2}\]

OpenStudy (anonymous):

In the first case, distributing all the factors gives \[\begin{align*} 2-5x^2&=A(1+x)^2+B(2-x)(1+x)+C(2-x)\\ &=(A-B)x^2+(2A+B-C)x+A+2B+2C \end{align*}\] In the second, \[\begin{align*} 2-5x^2&=A(1+x^2)+(Bx+C)(2-x)\\ &=(A-B)x^2+(2B-C)x+A+2C \end{align*}\]